A wave on a string is described by the relation y =0.15 sin(35t -0.025x), where t is measured

in seconds and x in meters. Find the amplitude, frequency, wavelength, and speed of this wave.


1
Expert's answer
2021-01-22T15:04:53-0500

Amplitude "0.15(m)"


Frequency "\nu=\frac{\omega}{2\pi}=\frac{35}{2\cdot3.14}=5.57(Hz)"


Wavelength "\lambda=\frac{2\pi}{k}=\frac{2\cdot3.14}{0.025}=251.3(m)"


Phase speed "v=\frac{\omega}{k}=\frac{35}{0.025}=1400(m/s)" or "v=\lambda\nu=251.3\cdot5.57=1400(m/s)"